OpenStudy (anonymous):

can anyone tell me if im right???? simplify 4x^3+7x^3 i got 11x^3

OpenStudy (lgbasallote):

riiiiight

OpenStudy (anonymous):

thank you!

Directrix (directrix):

4x^3+7x^3 = x^3 ( 4 + 7) = x^3 (11) = 11 x^3 YES

OpenStudy (anonymous):

yup! gj! you just add the constants in front of the x^3